One grade of oil has 0.50% of an additive and a higher grade has 0.75% of the additive. How many liters of each must be used to have 1000 L of a mixture with 0.65% of the aadditive? Please help solve. Thank You!
Answer by checkley77(12844)

.75X+.50(1,000-X)=.65*1,000
.75X+500-.50X=650
.25X=650-500
.25X=150
X=150/.25
X=600 AMOUNT OF 75% ADDITIVE IS USED.
1,000-600=400 AMOUNT OF 50% ADDITIVE IS USED.
PROOF:
.75*600+.50*400=.65*1,000
450+200=650
650=650